# EvAlignUX: Advancing UX Research through LLM-Supported Exploration of Evaluation Metrics

### v1

@@ 3.3. Technical Evaluation
We conducted a small-scale evaluation to compare EvAlignUX ’s understanding of research metrics with human understanding and assess the quality of its metric suggestions based on a given narration.
3.3.1. Evaluating LLM’s Ability to Extract the Metrics
@@
To assess GPT’s capability in identifying evaluation metrics in research papers from a predefined list of candidate metrics, we adopted a similar strategy as (Pang et al., 2024) and randomly selected 15 papers from our initial pool of research articles. The evaluation metrics used in these papers were extracted using our GPT prompt. One author reviewed each identified metric and rated it from three aspects: whether the metric was actually measured, whether the metric was in the candidate list and whether the reasoning or citation provided makes sense. A second author verified these assessments. A metric was considered accurate if it fulfils requirement one and two. Common sources of inaccuracy included:
The metric identified was not actually measured in the evaluation part of the paper (i.e. hallucination (Ji et al., 2023)). There were three cases for this: a) GPT inferred metrics based on descriptions from the paper, but the metrics were not explicitly measured in the evaluation process. b) GPT synthesized metrics from the results, but the metrics were not explicitly measured in the evaluation process and c) GPT making recommendations for a metric that should have been used but was not used, based on the context of the paper. The case for (a) and (b) was often that GPT picked out a statement or a conclusion from a qualitative result described in the paper. However, such qualitative results were synthesized from interviews or surveys, and therefore the identified metric was not explicitly measured in the paper. For example, GPT identified ”attitude towards the agent” as an evaluation metric from (Wambsganss et al., 2020), quoting ”students made positive comments about the interaction with CAs related to the perceived level of enjoyment and interaction.” While attitude towards the agent was not explicitly measured in the paper, it was described in the findings and was thus misidentified. For (c), an example is GPT identifying ”privacy concern” as a metric from (Lee et al., 2021), quoting ”because chatbot use can result in users’ self-disclosure of sensitive topics” from the discussion section. Although it makes sense to measure ”privacy concern” based on the context of potential self-disclosure of sensitive information, the said metric was not explicitly measure in the paper so it was marked as inaccurate. The metric identified was outside the list of candidate metrics. There were two cases for this: a) GPT reworded the metric in the candidate list to express the same meaning, and b) GPT made up a metric that was not included in the candidate list. For (a), an example was GPT identifying ”utterance length” from (Thomas et al., 2020), while the actual metric in the candidate list is worded as ”length of utterance”. Regarding (b), ”transparency” was identified as a metric from (Pecune et al., 2019), while ”transparency” is not part of the candidate list.
@@
We found that out of the 102 metrics identified in the 15 papers, 94% (96/102) were indeed measured, 92% (94/102) were in the candidate list and 88% were accurate (90/102), as there were two overlapping metrics across requirement one and two. 100% (102/102) of the reasoning and citations made sense. Overall, the results suggest that with our prompt, GPT’s ability to identify metrics is largely reliable.
3.3.2. Evaluation of Metric Recommendation
@@
We evaluated our system using a sample of 40 papers as ground truth, including papers in AI, wearable technology, VR/AR, conversational agents, etc. Our system is supported by GPT-4o, and we also benchmarked our system against the GPT-4o LLM model to recommend UX metrics using precision, recall, and F1 score (Tamm et al., 2021; Fayyaz et al., 2020). The baseline GPT-4o and our system were given the same set of inputs, including the system description and the corresponding list of indexes. Then, each sample was processed three times by both our system and the baseline to generate a list of metrics. The results obtained from both the systems were generalized to reflect the naming standards used in test dataset. To perform the final comparision, we calculated the precision, recall, and f1 score for each subsequent run, and the final metric scores obtained were averaged to obtain the overall mean scores.
@@
Table 1 demonstrates that our system significantly outperforms the baseline GPT-4o, with a 6% increase in precision, a 10.1% increase in recall, and a 7.4% increase in F1 score. We further conducted a pair-wise t-test by processing each sample 10 times on both the systems, to compare the metric mean between EvAlignUX and baseline (). The t-test demonstrate that their is there is a statistically significant difference between the performance (precision, recall, and F1 score) of the two models on the same dataset. Additionally, we observed that our system recommends more relevant UX metrics for a given system description than the baseline GPT-4o. For instance, when evaluating Tan et al.’s work (Tan et al., 2024), our system recommends using Perceived Naturalness to measure how naturally the AI-generated visuals integrate with the audiobook narrative. However, when evaluating the baseline GPT-4o, none of the generated metrics explicitly address the naturalness of the integration between visuals and audio. Furthermore, when testing the system on Wang et al (Wang et al., 2024), our system recommends evaluating Emotion Awareness to monitor the system’s ability to recognize and appropriately respond to users’ different emotional states, which is also measured in the research However, none of the metrics recommended by the baseline model address the evaluation of the varied emotional states of the user. Additionally, when using (Sun et al., 2024), the system recommends measuring Perceived Intelligence to assess how users perceive the system in generating coherent motions, which the baseline model does not explicitly recommend. The results indicate that our system outperforms the baseline GPT-4o in recommending UX metrics to users.
@@ 4. User Study
In this section, we outline how we evaluate EvAlignUX ’s user experience, and we aim to explore the following key research questions.
RQ1: How effectively does the use of EvAlignUX impact researchers in developing their UX evaluation plans?
RQ2: How do users leverage the EvAlignUX features to improve thought processes of developing UX evaluation plans?
RQ3: How does the users’ background impact their use of EvAlignUX ?
@@ 4.1. Participant
Our study was approved by the university’s IRB. We recruited 19 participants from social media. All participants reported they had extensive experience with HCI/UX research, 11 participants (57.9%) have 3-4 years experience, 7 participants (36.8%) have 1-2 years experience, only one participant has less than one year UX experience. Regarding participant’s AI usage in research study, the majority of participants used AI to help with simple tasks, such as coding & writing (17, 89.5%), literature review (7, 36.8%), and data analysis (11, 57.9%),. Notably, we found that few participants used AI for tasks that require rigor and creativity, including research ideaiton (7, 36.8%), design prototyping (7, 36.8%), research evaluation plan development (3, 15.8%), and usability testing (1, 5.3%). The background information of participants is shown in Appendix 6. Each user study lasted 1.5 to 2 hours. Participants received 20 USD per hour as compensation in the user study.
@@ 4.2. Procedure
Figure 6 demonstrates the study procedure. ① The study began with the background collection, including their UX experience levels. Researchers then introduced the study context and explained the tasks. Then, they used EvAlignUX ’s landing page to spend about 10 minutes entering their project description, evaluation plans, and expected outcomes. After that, they completed an initial survey that capture their perceptions of the proposed project plans (survey 1, initial plan). Then, participants used the Zoom Whiteboard to draw their thought processes for developing the UX evaluation plan and think-aloud to ensure their ideas were communicated clearly.
In EvAlignUX , participants were guided through the system features and asked to refine their project plans using all available features. After experimenting with the features and achieving satisfactory results, participants used the ”Export Data” button to review, regenerate, and re-edit their project plans. Upon finalizing their plans, they completed exit surveys (survey 2, revised plan; survey 3, EvAlignUX ) to evaluate their perceptions of the generated plans and their experience with the system. During the EvAlignUX process, the think-aloud method was employed to observe and record participants’ interactions and thoughts. Next, participants used Zoom Whiteboard to revise their thought process for developing the UX evaluation plan if they felt there were any changes. Finally, participants were interviewed to delve deeper into their overall experiences and gather additional insights. This step provided a comprehensive understanding of user perceptions and allowed for collecting feedback on both the proposed plans and the system features.
@@ 4.3.1. System Logs
For system log data, we firstly collected user’s input to the system, including project description (i.e. What are the main objectives of your project?), UX evaluation plans (i.e. How do you plan to evaluate the UX of your project?), and expected outcomes (i.e. What outcomes do you expect from this evaluation?) as well as their updated ones. Additionally, we logged their usage behavior data, including (1) the number of times a user regenerate indexes, (2) the total number of metrics selected, (3) the number of metrics added through graph, (4) the number of outcomes clicked including both selected and deselected, and (5) the number of risks clicked. We also logged their revised system description, UX evaluation plans, and expected outcomes.
@@ 4.3.2. Thought Process Diagrams
We collected user’s thought process visualization for developing a UX evaluation plan using Zoom Whiteboard before and after using the EvAlignUX to see how AI-based influenced their thinking.
@@ 4.3.4. Think Aloud
We collected their think-aloud while using the system features as well as their thought process diagrams. For example, we asked to provide us with reasoning behind their decision making process.
@@ 4.3.5. Interview
Finally, we conducted a short follow-up semi-structured interview. We asked questions in three main categories with follow-up questions based on the interviewee’s responses.
Perceptions of features and their role in decision making for proposing a UX plan: Which feature is the most impressive? Which feature is the least impressive? How do these features support your decision making?
@@ 4.4. Data Analysis
We analyze each participant’s ① thought process diagrams, ② quantitative data from surveys and logs, and ③ user decision making reasoning through think-aloud, interview transcripts, and UX research plans.
@@ 4.4.1. Thought Process Diagrams Analysis
We standardized thought process diagrams drawn by participants by coding each illustrated step. Modifications, such as envisioning AI assistance or adding new steps, were identified. These changes were further analyzed by examining their diagrams with the think-aloud data. We summarized the benefits participants identified where AI could assist in proposing a UX research plan.
@@ 4.4.2. Quantitative Data Analysis
We analyzed participants’ responses to the two post-task surveys. For all these measures, we conducted a Shapiro-Wilk test (González-Estrada and Cosmes, 2019) and found that all data were not normally distributed. Given the nature of our data, we applied Kendall’s tau correlation to explore relationships between user experience levels, system usage, and perceptions of changes in plans and system features(Bolboaca and Jäntschi, 2006; Kendall, 1938). We also reported effect sizes using standardized methodologies (Dixon and Massey Jr, 1951). All statistical analyses were conducted using Python 3.12.
@@ 4.4.3. Behavioral Analysis through Think-Aloud and Interviews
A thematic analysis (Braun and Clarke, 2012) was performed on data from think-aloud sessions and interviews. Two researchers collaboratively identified common themes through an iterative process. This analysis helped map how EvAlignUX features address identified challenges from the thought processes and outline the pros and cons of these features.
@@ 5. Findings
This section presents findings on how participants’ perceptions towards their proposed UX evaluation plans changed through the use of EvAlignUX (RQ1) in 5.1, their feature usages and perceptions (RQ2) in 5.2, and how these perceptions varied based on their UX research background (RQ2) in 5.3. Participant quotes are indicated by quotation marks, while system features are formatted in code style.
5.1. EvAlignUX Significantly Improving the UX Evaluation Plans (RQ1)
@@
In this section, we compared participants’ ratings of their original and revised UX plans across 11 measurement criteria. Of these, 9 measurements showed significant improvement in the revised plans compared to the original. Five participants fully accepted the system-generated evaluation plan without modification. On average, 60.1% (SD=0.33) of the generated UX evaluation plan and 59.6% (SD=0.38) of the generated UX outcome were incorporated into their final evaluation plans. The results are shown in Figure 7, with detailed information available in Appendix Table 7.
@@
Overall, the revised plans were more specific. The increase scores of Concreteness (W = 24.0, = .032∗, ), Completeness (W = 16.0, = .004∗∗, ), Rigor (W = 15.0, = .009∗∗, ), and Quality (W = 11.0, = 0.004∗∗, ) suggest that participants added more detailed information to their revised plan. Several participants only had a vague idea about which variables to measure (e.g., P7, P14, see Table 2). With EvAlignUX , they noted how it helped them improve the details in the UX plans. For example, P14 later strategically specified six measurements by questionnaires and system logs, with each in great detail; and P7 copied a holistic UX evaluation plan using a survey, system logs, and interview protocols, each with specific details added.
@@
Participants not only added more details to their initial plans, but they also perceived their plan more concise and relevant since Clarity (W = 4.0, = 0.005∗∗, ), and Efficiency in achieving research goals (W = 12.0, = .012∗, ) also demonstrated a considerable improvement. Participants made meaningful revisions for clarity. For instance, P5 removed texts related to assessment from their initial system description, as they found it more suitable as a part of the evaluation plan instead of putting it into the system narration. Also, several participants praised the "Metrics list" feature for clarifying metric relevance, noting it showed ”how this metric can be relevant in the context (P3)” and ”encouraged the reasoning of using this metric (P17)”. P8, during a think-aloud session, explained, ”I just selected ‘information retention’ because my project is about memory, and the definition reminds me of the importance of evaluating how well my participants retain information to measure the effectiveness of counseling sessions in my project.”
@@
Participants also felt more confident and believed their plans is feasible as Feasibility (W = 16.5, = 0.035∗, ) and Self-Confidence (W = 18.0, = .014∗, ) showed significantly increases. For example, participants felt they could know how to cite (P12, P15), ”measure (P1, P3, 14)”, ”design questions (P1, P9, P17)”, “learn previous findings (P8, P2)” for metrics using the metric feature, making the UX plan tangible, and even “how to collect the data specifically” (P13, see Table 2). They also gained confidence, especially when noticing prior work used the same metric selections. As P3 expressed, ”I’m really happy someone else also uses ’improvement in social skills’ in this context.” The validation further reinforced their self-confidence in the chosen metrics and plans.
Participants also significantly revised their plans to include more risk considerations (W = 10.0, = 0.007∗∗, ). For example, building upon original UX outcomes, P11 expanded, ”There alongside the expectations the research outcome may not align with education effect the system may not be generalizable. However, these may could be partially reflected in the objective metrics and be acknowledged in the limitations.” P11 took inspiration from concerns about generalizability from the system’s output (see Table 2, where ”Generalization Failure” was mentioned, stating, ”Thinking about the potential biases made me rethink how I approach user data to ensure that I don’t overlook important issues.”
However, there was no significant changes found in ”Novelty” (W = 22.0, = 0.083, ) and UX plans’ ”Relevancy” to research questions. Several participants felt the system was hard to change. P4 noted, ”The core idea is already set,” while P6 said, ”It’s challenging to introduce major changes.” Participants also reflected that ”It is hard to justify the novelty without referring to the actual paper and verify it on my own” (P4).

### v2

@@ 3.3. Technical Evaluation
We conducted a small-scale evaluation to compare EvAlignUX ’s understanding of evaluation metrics with human understanding and to assess the quality of its metric suggestions based on a given narrative. Figure 6 illustrates the process used to extract UX metrics applied in HCI publications and the method used to evaluate the accuracy of metric recommendations with EvAlignUX .
The graph shows a two-stage evaluation workflow: the top section illustrates the process of extracting UX evaluation metrics from full-text research papers from conferences like CHI, CSCW, DIS, and UIST are analyzed, metrics are automatically identified, and then human experts validate these extracted metrics to establish a ground truth set; the bottom section demonstrates how these ground truth metrics are used to evaluate recommendation tools, specifically comparing metrics suggested by a baseline model (GPT-4o) and our system (EvalignUX) using only paper abstracts and inferred indexes as inputs, with both recommendations assessed against the ground truth metrics using precision, recall, and F1 scores.
3.3.1. Evaluating Metrics Extraction
@@
To assess GPT’s capability in extracting evaluation metrics in research papers from a predefined list of candidate metrics, we adopted a similar strategy as in (Pang et al., 2024) and randomly selected 15 papers from our initial pool of research articles. The evaluation metrics used in these papers were extracted using our GPT prompt. One author reviewed each identified metric and rated it from three aspects: whether the metric was actually measured, whether the metric was in the candidate list, and whether the reasoning or citation provided makes sense. A second author verified these assessments. A metric was considered accurate if it fulfilled the above requirements.
@@
We found that out of the 102 metrics identified in the 15 papers, 94% (96/102) were indeed measured, 92% (94/102) were in the candidate list, and 88% were accurate (90/102), noting that two metrics overlapped between the first and second criteria. 100% (102/102) of the reasoning and citations made sense. Overall, the results suggest that with our prompt, GPT’s ability to identify metrics is largely reliable.
A common source of hallucination is that the metric extracted was not actually measured in the evaluation section of the paper (i.e., hallucination (Ji et al., 2023)). More specifically, there were three types of hallucination: a) GPT inferred metrics based on descriptions from the paper, but these metrics were not explicitly measured in the evaluation process; b) GPT synthesized metrics from the results, but these metrics were not explicitly measured in the evaluation process; and c) GPT made recommendations for a metric that should have been used but was not actually used, based on the context of the paper. The case for (a) and (b) was often that GPT often selected a statement or conclusion from a qualitative result described in the paper. However, such qualitative results were synthesized from interviews or surveys, and therefore the identified metric was not explicitly measured in the paper. For example, GPT identified ”attitude towards the agent” as an evaluation metric from (Wambsganss et al., 2020), quoting ”students made positive comments about the interaction with CAs related to the perceived level of enjoyment and interaction.” While attitude towards the agent was not explicitly measured in the paper, it was described in the findings and was thus misidentified. For (c), an example is GPT identifying ”privacy concern” as a metric from (Lee et al., 2021), quoting ”because chatbot use can result in users’ self-disclosure of sensitive topics” from the discussion section. Although it makes sense to measure ”privacy concern” based on the context of potential self-disclosure of sensitive information, this metric was not explicitly measured in the paper, so it was marked as inaccurate.
It was possible that the extracted metric was not in the list of candidate metrics. There were two reasons: (a) GPT reworded the metric in the candidate list while preserving its meaning, and (b) GPT invented a metric that was not included in the candidate list. For (a), one example is GPT identifying ”utterance length” from (Thomas et al., 2020), while the candidate list actually uses ”length of utterance”. Regarding (b), GPT identified ”transparency” as a metric from (Pecune et al., 2019), even though ”transparency” is not part of the candidate list.
Additionally, we observed two scenarios in which certain metrics were presented in the papers but not extracted by the model. First, as noted in (Wang et al., 2020b), GPT struggled to accurately extract metrics presented in tables or figures and omitted metrics such as ”fear of being judged” and ”usefulness.” Second, GPT omitted metrics that were described as ’measured’ in the original text but were not explicitly reported in the final results. For example, in (Kim et al., 2019), ”usage motivation” was ”measured” in the questionnaires but not reported. Similarly, in (Wambsganss et al., 2020), ”technology acceptance” was measured but not reported; instead, ”intention to use a conversational agent” was reported.
3.3.2. Evaluating Metrics Recommendation
We evaluated the performance of the metrics recommendation using a sample of 40 papers, including papers from CHI, CSCW, DIS, UIST, etc., in emerging technologies such as AI, wearable technology, VR/AR, conversational agents, and similar fields. The metrics for these papers were extracted using GPT-4o by providing the complete text of the papers to the LLM engine and they were manually verified by two authors to serve as the ground truth.
Our system is supported by GPT-4o, and we also compared our system against the same LLM engine to recommend UX metrics using precision, recall, and F1 score (Tamm et al., 2021; Fayyaz et al., 2020). The baseline GPT-4o and our system were given the same set of inputs, including the system description and the corresponding list of indexes. Unlike metrics extraction, where the entire text of the paper was provided, the evaluation against both EvAlignUX and the baseline LLM (GPT-4o) was conducted using only the system description as the input. This approach reflects the likelihood that EvAlignUX users would typically provide only a brief description of their system. Given the system description, a corresponding list of indexes was inferred using an LLM for each paper, as described in section 3.2.2. We then compared EvAlignUX and the baseline GPT-4o based on the number of metrics they recommended that were actually used in the papers, as determined by referring to the ground truth. For instance, Wan et al. (Wan et al., 2024) used “response quality”, “conversation quality”, “average response time”, and “agent behavior” to measure users’ satisfaction with NPC responses, evaluate the coherence and contextual relevance of NPC dialogues, determine the response latency, and assess the emotional authenticity of NPC interactions, respectively. Given the system description and identified indexes for this work, we assessed EvAlignUX ’s and the baseline’s ability to recommend relevant and coherent UX metrics.
@@ 3.3.2. Evaluating Metrics Recommendation
For implementation, we sampled each input three times using our system and the baseline to generate a list of metrics. The results obtained from both systems were generalized to reflect the naming standards used in the test dataset. To perform the final comparison, we calculated the precision, recall, and F1 score for each subsequent run, and the final metric scores obtained were averaged to obtain the overall mean scores.
@@
Table 2 demonstrates that our system significantly outperforms the baseline GPT-4o, with a 6% increase in precision, a 10.1% increase in recall, and a 7.4% increase in F1 score. We further conducted a pairwise t-test by processing each sample 10 times on both systems, to compare the metric mean between EvAlignUX and the baseline (). The t-test demonstrates that there is a statistically significant difference between the performance (precision, recall, and F1 score) of the two models on the same dataset. Additionally, we observed that our system recommends more relevant UX metrics for a given system description than the baseline GPT-4o. For instance, when evaluating Tan et al.’s work (Tan et al., 2024), our system recommends using Perceived Naturalness to measure how naturally the AI-generated visuals integrate with the audiobook narrative.
However, when evaluating the baseline GPT-4o, none of the generated metrics explicitly address the naturalness of the integration between visuals and audio. Furthermore, when testing the system on Wang et al. (Wang et al., 2024), our system recommends evaluating Emotion Awareness to monitor the system’s ability to recognize and appropriately respond to users’ different emotional states, which is also measured in the research. In contrast, none of the metrics recommended by the baseline model address the evaluation of the varied emotional states of the user. Additionally, when using the approach from (Sun et al., 2024), the system recommends measuring Perceived Intelligence to assess how users perceive the system’s ability to generate coherent motions, which the baseline model does not explicitly recommend. The results indicate that our system outperforms the baseline GPT-4o in recommending UX metrics to users.
@@ 4. User Study
In this section, we outline how we evaluate EvAlignUX ’s user experience and explore the following key research questions.
RQ1: How effectively does the use of EvAlignUX impact researchers when developing their UX evaluation plans?
RQ2: How do users leverage the EvAlignUX features to enhance their thought processes when developing UX evaluation plans?
RQ3: How do users’ backgrounds impact their use of EvAlignUX ?
@@ 4.1. Participant
Our study was approved by the university’s IRB. We recruited 19 participants from social media. All participants reported that they had extensive experience with HCI/UX research: 11 participants (57.9%) had 3–4 years of experience, 7 participants (36.8%) had 1–2 years of experience, and only one participant had less than one year of UX experience. Regarding participants’ use of AI in research studies, the majority used AI to help with simple tasks, such as coding and writing (17, 89.5%), literature review (7, 36.8%), and data analysis (11, 57.9%). Notably, we found that few participants used AI for tasks that require rigor and creativity, including research ideation (7, 36.8%), design prototyping (7, 36.8%), research evaluation plan development (3, 15.8%), and usability testing (1, 5.3%). The background information of participants is shown in Appendix 6. Each user study lasted 1.5 to 2 hours. Participants received 20 USD per hour as compensation for the study.
@@ 4.2. Procedure
Figure 7 demonstrates the study procedure. ① The study began with collecting background information, including participants’ UX experience levels. Researchers then introduced the study context and explained the tasks. Then, they used EvAlignUX ’s landing page to spend about 10 minutes entering their project descriptions, evaluation plans, and expected outcomes. After that, they completed an initial survey that captured their perceptions of the proposed project plans (survey 1, initial plan). Then, participants used the Zoom Whiteboard to draw their thought processes for developing the UX evaluation plan and engaged in a think-aloud to ensure their ideas were communicated clearly.
In EvAlignUX , participants were guided through the system features and asked to refine their project plans using all available features. After experimenting with the features and achieving satisfactory results, participants used the ”Export Data” button to review, regenerate, and re-edit their project plans. Upon finalizing their plans, they completed exit surveys (survey 2, revised plan; survey 3, EvAlignUX ) to evaluate their perceptions of the generated plans and their experience with the system. During the EvAlignUX process, the think-aloud method was employed to observe and record participants’ interactions and thoughts. Next, participants used Zoom Whiteboard to revise their thought processes for developing the UX evaluation plan if they felt changes were needed. Finally, participants were interviewed to delve deeper into their overall experiences and gather additional insights. This step provided a comprehensive understanding of user perceptions and allowed the collection of feedback on both the proposed plans and the system features.
This graph illustrates the study procedure for our user study. At the top, a timeline outlines the steps participants followed during the study. The process begins with participants completing a pre-survey to provide their background information. Next, they enter their initial project plan into EvalignUX. After submitting their plan, participants complete Survey 1, which captures their initial perceptions and requires them to draw their current research thought processes without the aid of AI. Following this, researchers present and explain key features of EvalignUX, allowing participants to explore metrics and finalize their plans. Participants then complete Survey 2, documenting their revised plan, followed by Survey 3, which evaluates the EvalignUX system’s features. They are also asked to update their thought process diagram using AI. The procedure concludes with an interview.
@@ 4.3.1. System Logs
For system log data, we first collected user’s input to the system, including the project description (i.e., What are the main objectives of your project?), UX evaluation plans (i.e. How do you plan to evaluate the UX of your project?), and expected outcomes (i.e. What outcomes do you expect from this evaluation?) as well as any updated ones. Additionally, we logged their usage behavior data, including (1) the number of times a user regenerated indexes, (2) the total number of metrics selected, (3) the number of metrics added through the graph, (4) the number of outcomes clicked including both selected and deselected outcomes, and (5) the number of risks clicked. We also logged their revised system description, UX evaluation plans, and expected outcomes.
@@ 4.3.2. Thought Process Diagrams
We collected users’ thought process visualizations for developing a UX evaluation plan using Zoom Whiteboard before and after using the EvAlignUX to see how AI-based assistance influenced their thinking.
@@ 4.3.4. Think Aloud
We recorded participants’ think-aloud protocols while they used the system features as well as their thought process diagrams. For example, we asked them to provide the reasoning behind their decision-making process.
@@ 4.3.5. Interview
Finally, we conducted a short follow-up semi-structured interview. We asked questions in three main categories, supplemented by follow-up questions based on the interviewee’s responses.
Perceptions of features and their role in decision-making for proposing a UX plan: Which feature is the most impressive? Which feature is the least impressive? How do these features support your decision-making?
@@ 4.4. Data Analysis
We analyze each participant’s ① thought process diagrams, ② quantitative data from surveys and logs, and ③ user decision-making reasoning through think-aloud protocols, interview transcripts, and UX research plans.
@@ 4.4.1. Thought Process Diagrams Analysis
We standardized the thought process diagrams drawn by participants by coding each illustrated step. Modifications, such as envisioning AI assistance or adding new steps, were identified. These changes were further analyzed by examining the diagrams alongside the think-aloud data. We summarized the benefits participants identified where AI could assist in proposing a UX research plan.
@@ 4.4.2. Quantitative Data Analysis
We analyzed participants’ responses to the two post-task surveys. For all these measures, we conducted a Shapiro-Wilk test (González-Estrada and Cosmes, 2019) and found that the data were not normally distributed. Given the nature of our data, we applied Kendall’s tau correlation to explore relationships between user experience levels, system usage, and perceptions of changes in plans and system features (Bolboaca and Jäntschi, 2006; Kendall, 1938). We also reported effect sizes using standardized methodologies (Dixon and Massey Jr, 1951). All statistical analyses were conducted using Python 3.12.
@@ 4.4.3. Behavioral Analysis through Think-Aloud and Interviews
A thematic analysis (Braun and Clarke, 2012) was performed on data from think-aloud sessions and interviews. Two researchers collaboratively identified common themes through an iterative process. This analysis helped map how EvAlignUX features address the challenges stemming from participants’ thought processes while also outlining the pros and cons of these features.
@@ 5. Findings
This section presents findings on how participants’ perceptions of their proposed UX evaluation plans changed through the use of EvAlignUX (RQ1) in Section 5.1, their feature usages and perceptions (RQ2) in 5.2, and how these perceptions varied based on their UX research background (RQ3) in Section 5.3. Participant quotes are indicated by quotation marks, while system features are formatted in code style.
5.1. EvAlignUX Significantly Improves the UX Evaluation Plans (RQ1)
@@
In this section, we compared participants’ ratings of their original and revised UX plans across 11 measurement criteria. Of these, nine measurements showed significant improvement in the revised plans compared to the original. Five participants fully accepted the system-generated evaluation plan without modification. On average, 60.1% (SD = 0.33) of the generated UX evaluation plan and 59.6% (SD = 0.38) of the generated UX outcome were incorporated into their final evaluation plans. The results are shown in Figure 8, with detailed information available in Appendix Table 7.
The image presents a bar chart comparing participants’ ratings on their proposed UX plans before (top bar of each item) and after (bottom bar of each item) using EvalignUX. Each item is evaluated across several dimensions, including “Self-Confidence”, “Trust”, “Considering Ethical Issues”, “Methodological Rigor”, “Feasibility”, “Quality”, “Concreteness”, “Clarity”, “Completeness”, “Efficiency”, “Relevance”, and “Novelty”. The bars are color-coded to represent varying degrees of rating: from very low to very high. Each item shows a noticeable shift in participant ratings, with the ratings of after using EvalignUX generally leaning towards higher values, indicated by the darker shades of blue.
@@
Overall, the revised plans were more specific. The increased scores in Concreteness (W = 24.0, = .032∗, ), Completeness (W = 16.0, = .004∗∗, ), Rigor (W = 15.0, = .009∗∗, ), and Quality (W = 11.0, = 0.004∗∗, ) suggest that participants added more detailed information to their revised plan. Several participants only had a vague idea about which variables to measure (e.g., P7, P14, see Table 2). With EvAlignUX , they noted how it helped them improve the details in the UX plans. For example, P14 later strategically specified six measurements by questionnaires and system logs, with each in great detail; and P7 copied a holistic UX evaluation plan using a survey, system logs, and interview protocols, each with specific details added.
@@
Participants not only added more details to their initial plans but also perceived them as more concise and relevant, as both Clarity (W = 4.0, = 0.005∗∗, ), and Efficiency in achieving research goals (W = 12.0, = .012∗, ) also demonstrated a considerable improvement. Participants made meaningful revisions for clarity. For instance, P5 removed text related to assessment from their initial system description, as they found it more suitable as a part of the evaluation plan instead of putting it into the system narration. Also, several participants praised the "Metrics list" feature for clarifying metric relevance, noting it showed ”how this metric can be relevant in the context (P3)” and ”encouraged the reasoning of using this metric (P17)”. P8, during a think-aloud session, explained, ”I just selected ‘information retention’ because my project is about memory, and the definition reminds me of the importance of evaluating how well my participants retain information to measure the effectiveness of counseling sessions in my project.”
@@
Table 2 presents examples of participants’ initial and revised UX evaluation plans after using EvalignUX. Each row corresponds to either a “UX Plan” or “UX Outcome” as the input type from participants (i.e., P7, P14, P13, P11). The “Initial Plan” column summarizes the participant’s original approach to evaluating user experience. The “Revised Plan” column shows how these plans evolved, providing more specific methods with the aid of EvalignUX. Through these examples, Table 2 illustrates how participants refined their UX evaluation strategies to gather richer, more actionable insights.
Participants also felt more confident and believed their plans were feasible as Feasibility (W = 16.5, = 0.035∗, ) and Self-Confidence (W = 18.0, = .014∗, ) as indicated by significant increases. For example, participants felt they could know how to cite (P12, P15), ”measure (P1, P3, P14)”, ”design questions (P1, P9, P17)”, “learn previous findings (P8, P2)” for metrics using the metric feature, making the UX plan tangible, and even “how to collect the data specifically” (P13, see Table 2). They also gained confidence, especially when noticing prior work used the same metric selections. As P3 expressed, ”I’m really happy someone else also uses’ improvement in social skills in this context.” The validation further reinforced their self-confidence in the chosen metrics and plans.
Participants also significantly revised their plans to include more risk considerations (W = 10.0, = 0.007∗∗, ). For example, building upon their original UX outcomes, P11 expanded, ”There alongside the expectations the research outcome may not align with education effect the system may not be generalizable. However, these issues could be partially reflected in the objective metrics and be acknowledged in the limitations.” P11 took inspiration from concerns about generalizability from the system’s output (see Table 2, where ”Generalization Failure” was mentioned, stating, ”Thinking about the potential biases made me rethink how I approach user data to ensure that I don’t overlook important issues.”
However, there were no significant changes found in ”Novelty” (W = 22.0, = 0.083, ) or UX plans’ ”Relevancy” to research questions. Several participants felt that the system was hard to change. P4 noted, ”The core idea is already set,” while P6 said, ”It’s challenging to introduce major changes.” Participants also reflected that ”It is hard to justify the novelty without referring to the actual paper and verifying it on my own” (P4).